daboyjaylen:

If s = 1 over 4 unit and A = 16s2, what is the value of A, in square units? square unit(s).

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \ A = 16s^2 A = \[A = 16\left( \frac{ 1 }{ 4 } \right)^2 = ?\] \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

@Laylalyssa here

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa here \(\color{#0cbb34}{\text{End of Quote}}\) Not rly sure how to do this 1

jhonyy9:

just substitute in place of s the given value